Round tablets to the nearest 1/2 tablet. Round IV calculation to the nearest whole number. 1.An...

  1. Round tablets to the nearest 1/2 tablet. Round IV calculation to the nearest whole number.

  2. 1.An order reads: Give 600 mL over 5 hours at 20 gtt/mL. What is the flow rate in _______gtt/min?

(Remember to round your answer to a whole number)

  1. The nurse has a child with impetigo and receives orders to start the antibiotic cephalexin (Keflex)50 mg/kg/24 hours divided into three equal doses for a child that weighs 33 lbs. The medication is provided in 250-mg tablets. How many tablets will the nurse provide with each dose?_____tablet

  1. What is the flow rate per hour of 1000mL to run over 5 hours? _____mL/hr.  

Remember with IV's to round to the nearest whole number)

  1. The nurse is tallying intake for a client with congestive heart failure who is on a 1,000-mL fluid restriction. The client has had 1 cups of coffee, 1/2 cup of water, 2 oz juice, 30% of a 10 oz cup of soup, and half a pint of milk. How much additional fluid in milliliters can the client consume for the remainder of the day?

  1. A patient requires a 30% decrease in the dosage of his medication. His current dosage is 340 mg. What will his dosage be after the decrease?

  1. The order reads" diphenhydramine (Benedryl) 50mg. The bottle reads diphenhydramine 25mg/5mL. How many mL do you give?________mL

  1. If the daily dosage of celebrex is 200 mg and the medication is given qid, how many mg. do you give per dose?______mg.

  1. You have an order to administer 1000 mg of metformin and have 0.5g tablets available. How many tablets will you administer?
Answer #1

1)600 ml has to be administered over 5 hours.

Per ml = 20 gtt

So 600 ml = 600*20 = 12000 gtt

So 12000 gtt in 5 hours.

=12000 gtt in 300 minutes

Per minute = 12000/300= 40 gtt/min

So answer is 40 gtt/min.

Weight of the child = 33 lbs = 33*0.454= 14.98 kg

Total dose = 50 mg * 14.98 = 749.1 mg

It has to be given in 3 doses.

So per dose = 749.1/3 = 250 mg

So per dose 1 tablet should be given.

Flow rate = 1000 ml/5 hours

= 200 ml/hour

His dosage after the decrease= 340 - 30/100*340 = 238 mg

25 mg in 5 ml.

So 50 mg= 5/25 * 50 = 10 ml

Qid means 4 times a day.

So per dose = 200/4 =50 mg

0.5 gram =500 mg

So 1 tablet should be administered.